Employers often use standardized measures to gauge how likely it is that a new employee with little to no experience will succeed in their company. One such factor is intelligence, measured using the Intelligence Quotient (IQ). To show that this factor is related to job success, an organizational psychologist measures the IQ score and job performance (in units sold per day) in a sample of 10 new employees.

IQ Job Performance
100 17
115 37
108 32
98 15
120 45
147 57
132 46
85 55
105 28
110 38

(a) Convert the following data to ranks and then compute a Spearman correlation coefficient. (Round your answer to three decimal places.)


(b) Using a two-tailed test at a 0.05 level of significance, state the decision to retain or reject the null hypothesis.

Retain the null hypothesis.Reject the null hypothesis.
Answer #1

to rank the IQ and Tob peo formance first we reed Seperately. Hollowing table shows the ranks, IQ Job performance Rank (IQ) h1Q Job performance Bank (19) Romk Job Performance) diz Rank(IG) Rank (Tob Performance) -1 lo 17. C UỐI 37 08 32 9 o 15 147 57© Degrees of freedom of-n-2 =10- 28 [df=8 P-value two-tailed test is between 0.10 and 0.20 (.. By using Spearmans Rank corre